The driver driving a university bus starts from rest and accelerates uniformly for 10s to a velocity of 36km/h during a test driver after servicing it. He then runs at a constant velocity and finally brought to rest in 50m with a constant retardation. The total distance covered during the test drive is 500m.

(a) Sketch the velocity – time graph for this test drive of the bus.

(b) Find the value of;

(i) acceleration

(ii) retardation

(iii) total time taken



1
Expert's answer
2021-03-02T18:08:17-0500

a)



b) i)


"a=\frac{10}{10}=1\frac{m}{s^2}"

ii)


"10^2=2(50)a'\\a'=1\frac{m}{s^2}"

iii)


"T=10+\frac{500-50-0.5(10)(10)}{10}+\frac{2\cdot50}{10}=60\ s"